# Competition of Mechanisms: Tracing How Language Models Handle Facts and Counterfactuals

### v1

@@ Abstract
Interpretability research aims to bridge the gap between the empirical success and our scientific understanding of the inner workings of large language models (LLMs). However, most existing research in this area focused on analyzing a single mechanism, such as how models copy or recall factual knowledge. In this work, we propose a formulation of competition of mechanisms, which instead of individual mechanisms focuses on the interplay of multiple mechanisms, and traces how one of them becomes dominant in the final prediction. We uncover how and where the competition of mechanisms happens within LLMs using two interpretability methods, logit inspection and attention modification. Our findings show traces of the mechanisms and their competition across various model components, and reveal attention positions that effectively control the strength of certain mechanisms. [^1] Our code and data are at https://github.com/francescortu/Competition_of_Mechanisms.
@@ 1 Introduction
We propose a novel formulation of competition of mechanisms, which focuses on tracing each mechanism in the model, and understanding how one of them becomes dominant in the final prediction by winning the “competition”. Specifically, we build our work on two single mechanisms that are well-studied separately in literature: (1) the factual knowledge recall mechanism, which can be located in the MLP layers (Geva et al., 2021; Meng et al., 2022; Geva et al., 2023); and (2) the in-context adaptation to a counterfactual statement, which is enabled by the copy mechanism conducted by induction heads of attention layers Elhage et al. (2021); Olsson et al. (2022). Based on the latest tools to inspect each of these two mechanisms Nostalgebraist (2020); Wang et al. (2023); Geva et al. (2023), we then unfold how and where the competition of the two mechanisms happen, and how it leads to the overall success or failure of LLMs.
Technically, we deploy two main methods: logit inspection Nostalgebraist (2020); Geva et al. (2022) by projecting the outputs of each model components by an unembedding matrix, and attention modification Geva et al. (2023); Wang et al. (2023). Using these methods, we assess the contributions of various model components, both from a macroscopic view (e.g., each layer) as well as a microscopic view (e.g., attention heads), and identify critical positions and attention heads involved in the competition of the two mechanisms. Moreover, we locate a few localized positions of some attention head matrices that can significantly control the strength of the factual mechanism. We summarize our main findings as follows:
@@
    In early layers, the factual attribute is encoded in the subject position, while the counterfactual is in the attribute position (Sec. 6.1);
@@
    The attention blocks write most of the factual and counterfactual information to the last position (Sec. 6.2);
@@
    All the highly activated heads attend to the attribute position regardless of the specific type of information they promote. The factual information flows by penalizing the counterfactual attribute rather than promoting the factual one (Sec. 6.3);
@@
    We find that we can up-weight the value of a few, very localized values of the attention head matrix to strengthen factual mechanisms substantially (Sec. 6.4).
@@ 2 Related Work on Interpretability
One major type of work in interpretability has focused on understanding what has been encoded in the representations of deep learning models. (Alain and Bengio, 2016; Conneau et al., 2018; Hupkes et al., 2018; Hewitt and Liang, 2019; Tenney et al., 2019; Jiang et al., 2020; Elazar et al., 2021, inter alia). Example features of interest include part of speech Belinkov et al. (2017), verb tense Conneau et al. (2018), syntax Hewitt and Manning (2019), and factual knowledge Petroni et al. (2019).
Interpreting the Learned Mechanisms/Functions.
Beyond interpreting the representations in the hidden states of the black-box models, another research direction is to interpret the mechanisms or functions that the models have learned, giving rise to the field of mechanistic interpretability (Olah et al., 2020; Elhage et al., 2021; Olsson et al., 2022; Nanda et al., 2023, inter alia). Some example mechanisms decoded in recent work include: mathematical operations such as modular addition Nanda et al. (2023) and the greater-than operation (Hanna et al., 2023); natural language-related operations such as the copy mechanism achieved by induction heads in LLMs Olsson et al. (2022) and factual knowledge recall achieved by MLP layers (Geva et al., 2021; Meng et al., 2022; Geva et al., 2023), which we describe below.
@@ 3 Problem Setup
Next, we design the task to incorporate the competition of mechanisms as in Figure 1. Specifically, for each factual statement $\bm{f}:=(t_{1}^{f},\dots,t_{k}^{f})$ consisting of $k$ tokens (e.g., “iPhone was developed by Apple”), we compose a corresponding counterfactual statement $\bm{c}:=(t_{1}^{c},\dots,t_{k^{\prime}}^{c})$ (e.g., “iPhone was developed by Google”). Then, we compose a prompt connecting the two statements as “Redefine: $\bm{c}$. $\bm{f}_{1:k-1}$.”, such as “Redefine: iPhone was developed by Google. iPhone was developed by ___”.
@@ 4 Method and Background
To inspect the inner workings of the two mechanisms, we trace the residual stream Elhage et al. (2021), or logits of each component in the LLM. Given a text sequence of $k$ tokens, LLMs maps it into the residual stream, namely a matrix $\mathbf{x}\in\mathbb{R}^{d\times k}$, where $d$ is the dimension of the internal states of the model. We use the term $\mathbf{x}_{i}^{l}$ to specify the residual stream at position $i$ and layer $l$.
@@
Modifying or ablating the activation of a specific model component is also a strategy used to improve the understanding of the information flow within LLMs, including techniques such as causal tracing Meng et al. (2022) and attention knockout (Wang et al., 2023; Geva et al., 2023). In our work, we mainly modify the attention matrix ${A}^{hl}$ of a certain attention head of the attention layer $\mathbf{a}^{l}$. Formally, in the attention head $h$ at layer $l$, we can force-set the attention weight of a token $\mathbf{x}_{i}^{l}$ to one of its earlier tokens $\mathbf{x}_{j}^{l}$ to be $\alpha$:
${A}_{ij}^{hl}=\alpha,\quad\text{ where }j<i~{}.$ (5)
@@ 5 Experimental Setup
To compose the factual and counterfactual statements as introduced in Sec. 3, we adopt CounterFact (Meng et al., 2022), commonly used dataset to interpret models’ ability of factual knowledge recall. We select 10K data points by considering only examples where the attributes are represented by a single token and where the model completes the sentence correctly. See our data composition details in Appendix A.
Figure 2: Analysis of the key positions of the factual and counterfactual in GPT-2. This figure displays the logit values for two significant tokens across different positions and layers. The upper left panel shows the logit values for the factual token $t_{\mathrm{fact}}$. Logit($t_{\mathrm{fact}}$) is higher in the subject position in the initial layers and in the last position of the premise and second sentence in the final layers. The lower right panel shows the logit of the counterfactual token $t_{\mathrm{cofa}}$. Logit($t_{\mathrm{cofa}}$) is higher in the attribute position in the first layers and in the last position of the second sentence at the end of the network. The right panel shows the average logits of $t_{\mathrm{fact}}$ (red) and of $t_{\mathrm{cofa}}$ (blue) in the last sequence position. At the output of the network Logit($t_{\mathrm{cofa}}$) > Logit($t_{\mathrm{fact}}$) in 96% of the examples.
@@
We follow most existing interpretability studies (e.g., Meng et al., 2022; Wang et al., 2023; Conmy et al., 2023; Hanna et al., 2023) to use the autoregressive transformer, GPT-2 small (Radford et al., 2019) has 117M parameters, consisting of 12 layers with 12 self-attention heads each and a residual stream of 768 dimensions. We present our main results using GPT-2 small and also provide supplemental results of a larger model, Pythia-6.9B (Biderman et al., 2023) in Appendix C, to show the robustness of our findings across LLMs of different scale. In particular, since Pythia-6.9B has 32 layers with 32 self-attention heads each and a model dimension of 4096, results for this model reflect the consistency of our findings over a 30x increase in the number of parameters. We introduce our experimental details in Appendix B.
@@ 6 Results and Findings
In this section, we trace the competition of the mechanisms within the LLM via the two methods introduced in Sec. 4, i.e., inspecting the residual stream and intervening on the attention. We provide mechanistic analyses on five research questions in the following subsections:
@@
    Macroscopic view: In which layers do the competition of the mechanisms take place? (Sec. 6.1)
@@
    Intermediate view: How do we attribute the prediction to attention and MLP blocks? (Sec. 6.2)
@@
    Microscopic view: How do individual attention heads contribute to the prediction? (Sec. 6.3)
@@
    Intrinsic intervention: Can we modify the strength of a certain mechanism? (Sec. 6.4)
@@
    Behavioral intervention: What word choice varies the strength of the counterfactual mechanism in the given context? (Sec. 6.5)
6.1 Where Do the Two Mechanisms Take Place?
In the main model that we inspect, GPT-2, we find that it can usually identify the correct mechanism, the counterfactual mechanism, in 96% of the 10K test examples. This means that, in the last sequence position, at the output of the network, the counterfactual token, $t_{\mathrm{cofa}}$, gets most of the times a higher probability than $t_{\mathrm{fact}}$. In the following, we will inspect how the “winning” of the counterfactual mechanism happens across the layers of the LLM.
Methods.
We study how $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$ are encoded in the residual stream using the first method described previously in Sec. 4. Specifically, for a given position $i$ and a layer $l$, we project the embedding $\mathbf{x}_{i}^{l}$ (see Eq.(2)) to the vocabulary space using the last layer unembedding matrix $\mathbf{\tilde{x}}_{i}^{l}=W_{U}\mathbf{x}_{i}^{l}$. By varying $l$, we measure the value of the logit of $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$ as they evolve in the residual stream after the first attention block.
@@ Results.
The upper panel of Fig. 2-left shows the logit of the factual token, $t_{\mathrm{fact}}$. In the early layers, Logit($t_{\mathrm{fact}}$) is higher in the subject positions, and in late layers, its prevalence shifts to the last token of the sentences. A high logit value in the last position of late layers is expected, as this position is used to predict the subject attribute. Consistently with Geva et al. (2023), we also found that the factual attribute is written in the subject positions by the first MLP, which increases on average the value of $t_{\mathrm{fact}}$ from $3.3$ to $13.1$ in the premise and from $9.7$ to $13.5$ in the second sentence. The evolution of $t_{\mathrm{cofa}}$’s logits is shown in the lower panel. In this case, $t_{\mathrm{cofa}}$ is best encoded in the attribute position in early layers. Indeed, after the first attention block, the average logit value of $t_{\mathrm{cofa}}$ is $19.8$, and it is, in most of the examples, the highest in the $50k$-dimensional vocabulary of GPT-2.
Figure 3: Contributions of attention and MLP blocks to the logits of $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$ in the last sequence position. The figure shows the contribution of attention and MLP blocks to the logit difference $\Delta_{\mathrm{cofa}}=\text{Logit}(t_{\mathrm{cofa}})-\text{Logit}(t_{\mathrm{fact}})$ at $W_{U}\mathbf{a_{N}}^{l}$ (resp. $W_{U}\mathbf{m_{N}}^{l}$) for the input’s last position across layers. The attention blocks (left) contribute more significantly than the MLP blocks (right). Only the attention block at layer $11$ favors $t_{\mathrm{fact}}$.
As we proceed towards the output, the rank of $t_{\mathrm{cofa}}$ in the projected logit distribution remains very low: $t_{\mathrm{cofa}}$ is among the 20 most likely tokens in the first five layers and between the 20th and the 70th in the last part of the network (see Appendix, Fig. S5). Figure 2-right shows that after four attention blocks, the logit of the counterfactual attribute grows markedly in the last position of the second sentence, becoming higher than Logit($t_{\mathrm{cofa}}$) and reaching its peak value of 30 at layer 10. In the last layer, the logit of $t_{\mathrm{cofa}}$ slightly decreases but remains larger than that of $t_{\mathrm{fact}}$ in 96% of the examples.
The pattern we have described so far shows that the information about $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$ is brought to the last position after 6 or 7 layers (see Fig. 2-right). We have also seen that the most activated positions at early layers are the subjects for $t_{\mathrm{fact}}$ and the attribute for $t_{\mathrm{cofa}}$. The information flow between different sequence positions suggests a major role played by the attention mechanism in moving such information to the last position, as observed in Geva et al. (2023). We continue our analysis studying the role of the self-attention and MLP blocks in writing the information about $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$, and by understanding which, between subject and attribute positions, is the main source of factual information written to the last position in the final layers.
Figure 4: Attention pattern for relevant attention heads. Left: Direct contribution to $\Delta_{\mathrm{cofa}}$ of all heads in GPT-2. Heads favoring $t_{\mathrm{fact}}$ are colored in red, and those favoring $t_{\mathrm{cofa}}$ in blue. Right: attention scores of relevant heads for the last position. A large attention score in the attribute position is found in all highly activated heads.
6.2 Counterfactual Contributions of Attention and MLP Blocks
Methods.

### v2

@@ Abstract
Interpretability research aims to bridge the gap between empirical success and our scientific understanding of the inner workings of large language models (LLMs). However, most existing research focuses on analyzing a single mechanism, such as how models copy or recall factual knowledge. In this work, we propose a formulation of competition of mechanisms, which focuses on the interplay of multiple mechanisms instead of individual mechanisms and traces how one of them becomes dominant in the final prediction. We uncover how and where mechanisms compete within LLMs using two interpretability methods: logit inspection and attention modification. Our findings show traces of the mechanisms and their competition across various model components and reveal attention positions that effectively control the strength of certain mechanisms.111 Code: https://github.com/francescortu/comp-mech. Data: https://huggingface.co/datasets/francescortu/comp-mech.
@@ 1 Introduction
We propose a novel formulation of competition of mechanisms, which focuses on tracing each mechanism in the model and understanding how one of them becomes dominant in the final prediction by winning the “competition”. Specifically, we build our work on two single mechanisms that are well-studied separately in literature: (1) the factual knowledge recall mechanism, which can be located in the MLP layers (Geva et al., 2021; Meng et al., 2022; Geva et al., 2023); and (2) the in-context adaptation to a counterfactual statement, which is enabled by the copy mechanism conducted by induction heads of attention layers Elhage et al. (2021); Olsson et al. (2022). Based on the latest tools to inspect each of these two mechanisms Nostalgebraist (2020); Wang et al. (2023); Geva et al. (2023), we then unfold how and where the competition of the two mechanisms happen, and how it leads to the overall success or failure of LLMs.
Technically, we deploy two main methods: logit inspection Nostalgebraist (2020); Geva et al. (2022) by projecting the outputs of each model component by an unembedding matrix, and attention modification Geva et al. (2023); Wang et al. (2023). Using these methods, we assess the contributions of various model components, both from a macroscopic view (e.g., each layer) and a microscopic view (e.g., attention heads), and identify critical positions and attention heads involved in the competition of the two mechanisms. Moreover, we locate a few localized positions of some attention head matrices that can significantly control the strength of the factual mechanism. We summarize our main findings as follows:
@@
    In early layers, the factual attribute is encoded in the subject position, while the counterfactual is in the attribute position (Section 6.1);
@@
    The attention blocks write most of the factual and counterfactual information to the last position (Section 6.2);
@@
    All the highly activated heads attend to the attribute position regardless of the specific type of information they promote. The factual information flows by penalizing the counterfactual attribute rather than promoting the factual one (Section 6.3);
@@
    We find that we can up-weight the value of a few very localized values of the attention head matrix to strengthen factual mechanisms substantially (Section 6.4).
@@ 2 Related Work on Interpretability
One major type of work in interpretability has focused on understanding what has been encoded in the representations of deep learning models. (Alain and Bengio, 2016; Conneau et al., 2018; Hupkes et al., 2018; Hewitt and Liang, 2019; Tenney et al., 2019; Jiang et al., 2020; Elazar et al., 2021, inter alia), or by geometric methods Doimo et al. (2020); Valeriani et al. (2024); Park et al. (2024); Cheng et al. (2024). Example features of interest include part of speech Belinkov et al. (2017), verb tense Conneau et al. (2018), syntax Hewitt and Manning (2019), and factual knowledge Petroni et al. (2019).
Interpreting the Mechanisms/Functions.
Beyond interpreting the representations in the hidden states of the black-box models, another research direction is to interpret the mechanisms or functions that the models have learned, giving rise to the field of mechanistic interpretability (Olah et al., 2020; Elhage et al., 2021; Olsson et al., 2022; Nanda et al., 2023, inter alia). Some example mechanisms decoded in recent work include mathematical operations such as modular addition Nanda et al. (2023) and the greater-than operation (Hanna et al., 2023); natural language-related operations such as the copy mechanism achieved by induction heads in LLMs Olsson et al. (2022) and factual knowledge recall achieved by MLP layers (Geva et al., 2021; Meng et al., 2022; Geva et al., 2023), which we describe below.
@@ 3 Problem Setup
Next, we design the task to incorporate the competition of mechanisms as in Figure 1. Specifically, for each factual statement $\bm{f}:=(t_{1}^{f},\dots,t_{k}^{f})$ consisting of $k$ tokens (e.g., “iPhone was developed by Apple.”), we compose a corresponding counterfactual statement $\bm{c}:=(t_{1}^{c},\dots,t_{k^{\prime}}^{c})$ (e.g., “iPhone was developed by Google.”). Then, we compose a prompt connecting the two statements as “Redefine: $\bm{c}$. $\bm{f}_{1:k-1}$.”, such as “Redefine: iPhone was developed by Google. iPhone was developed by ___”.
@@ 4 Method and Background
To inspect the inner workings of the two mechanisms, we trace the residual stream Elhage et al. (2021), or logits of each component in the LLM. Given a text sequence of $k$ tokens, LLMs map it into the residual stream, namely a matrix $\mathbf{x}\in\mathbb{R}^{d\times k}$, where $d$ is the dimension of the internal states of the model. We use the term $\mathbf{x}_{i}^{l}$ to specify the residual stream at position $i$ and layer $l$.
@@
Known as the Logit Lens (Nostalgebraist, 2020), this method is broadly adopted due to its consistent success in yielding interpretable results, demonstrating its effectiveness through broad empirical usage. However, it is important to note that it can occasionally fail to reflect the actual importance of vocabulary items, especially in the early layers of the network (Belrose et al., 2023).
@@
Modifying or ablating the activation of a specific model component is also a strategy used to improve the understanding of the information flow within LLMs, including techniques such as causal tracing Meng et al. (2022) and attention knockout (Wang et al., 2023; Geva et al., 2023).
In our work, we focus on modifying a small number of entries in the attention matrix. Namely, in the attention matrix ${A}^{hl}$ of the $h$-th head of the $l$-th attention layer $\mathbf{a}^{l}$, we focus on a certain entry, e.g., at the $(i,j)$ position, where $j<i$, which is the attention value of the token $\mathbf{x}_{i}^{l}$ attending to one of its earlier tokens $\mathbf{x}_{j}^{l}$. Following recent work Yu et al. (2023) , the modification is after the softmax layer, so the other attention values of the matrix stay unchanged. For the target entry ${A}_{ij}^{hl}$, we scale it up by a multiplier of $\alpha$:
${A}_{ij}^{hl}\leftarrow\alpha\cdot{A}_{ij}^{hl},\quad\text{ where }j<i~{}.$ (5)
@@ 5 Experimental Setup
To compose the factual and counterfactual statements as introduced in Section 3, we adopt CounterFact [^2] https://rome.baulab.info/data/ (Meng et al., 2022), commonly used dataset to interpret models’ ability of factual knowledge recall. We select 10K data points by considering only examples where the attributes are represented by a single token and where the model completes the sentence in a factually accurate manner.
Each instance of CounterFact expresses a relation $r$ between a subject $s$ and an attribute $a$: $(s,r,a)$. For example, in the sentence “iPhone was developed by Apple”, $s=$ “iPhone”, $r=$ “was developed by”, $a=$“Apple”. Moreover, each $(s,r)$ instance is provided two values of the attribute $a$, namely a factual token $t_{\mathrm{fact}}$, and a counterfactual token $t_{\mathrm{cofa}}$, representing a false fact.
[…]
(a) The logit values for the factual token $t_{\mathrm{fact}}$ (blue), and counterfactual token $t_{\mathrm{cofa}}$ (red).
(b) The average logits of $t_{\mathrm{fact}}$ (blue) and $t_{\mathrm{cofa}}$ (red) in the last token position.
Figure 2: Logits of the factual token $t_{\mathrm{fact}}$ and counterfactual token $t_{\mathrm{cofa}}$ across different positions and layers in GPT-2. The logit of $t_{\mathrm{fact}}$ is higher in the subject position in the initial layers and in the last position of the premise and second sentence in the final layers. The logit of $t_{\mathrm{cofa}}$ is higher in the attribute position in the first layers and in the last position of the second sentence at the end of the network.
@@
For this work, we first choose the GPT-2 small (Radford et al., 2019) model as it is the most commonly used one in previous interpretability studies (e.g., Meng et al., 2022; Wang et al., 2023; Conmy et al., 2023; Hanna et al., 2023). Aligning the same model with those studies can communicate the findings of this work better in the context of existing literature. Then, in addition to GPT-2, we check the generalizability of our work by provide supplemental results of Pythia-6.9B (Biderman et al., 2023) in Appendix A, to show the robustness of our findings across the two LLMs of different architectures and scales. In this way, having similar results across the two very diverse models makes the finding stronger than existing studies, most of which are only on GPT-2.
Implementation Details
As for experimental details, GPT-2 small has 117M parameters, consisting of 12 layers with 12 self-attention heads each and a residual stream of 768 dimensions. Pythia-6.9B has 32 layers with 32 self-attention heads each and a model dimension of 4,096, with a 30x increase in the number of parameters. For all our experiments, we deploy the pre-trained models from the Huggingface Hub Wolf et al. (2020), and inspect the residual streams by the LogitLens tool in the TransformerLens library (Nanda and Bloom, 2022).
@@ 6 Results and Findings
In this section, we trace the competition of the mechanisms within the LLM via the two methods introduced in Section 4, i.e., inspecting the residual stream and intervening on the attention. We provide mechanistic analyses on five research questions in the following subsections:
@@
    Macroscopic view: Which layers and token positions contribute to the two mechanisms? (Section 6.1)
@@
    Intermediate view: How do we attribute the prediction to attention and MLP blocks? (Section 6.2)
@@
    Microscopic view: How do individual attention heads contribute to the prediction? (Section 6.3)
@@
    Intrinsic intervention: Can we edit the model activations to modify the strength of a certain mechanism? (Section 6.4)
@@
    Behavioral analysis: What word choice varies the strength of the counterfactual mechanism in the given context? (Section 6.5)
6.1 Macroscopic Inspection across Layers and Token Positions
In the main model that we inspect, GPT-2, we find that it can usually identify the counterfactual mechanism in 96% of the 10K test examples. This means that, in the last sequence position, at the output of the network, the counterfactual token, $t_{\mathrm{cofa}}$, gets most of the times a higher probability than $t_{\mathrm{fact}}$. In the following, we will inspect how the “winning” of the counterfactual mechanism happens across the layers of the LLM.
Method.
We study how $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$ are encoded in the residual stream using the logit inspection method described in Section 4. Specifically, for a given token position $i$ and a layer $l$, we project the embedding $\mathbf{x}_{i}^{l}$, i.e., the residual stream in Eq. 2, to the vocabulary space by $\mathbf{\tilde{x}}_{i}^{l}=W_{U}\cdot\mathrm{norm}(\mathbf{x}_{i}^{l}$), where $W_{U}$ is the unembedding matrix and $\mathrm{norm}$ is the normalization of the last layer of the model. By varying $l$, we measure the values of the logits of $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$ as they evolve in the residual stream after the first attention block.
(a) Logit difference $\Delta_{\mathrm{cofa}}$ of the last token position after the attention block in each layer of GPT-2.
(b) Logit difference $\Delta_{\mathrm{cofa}}$ of the last token position after the MLP block in each layer of GPT-2.
Figure 3: Contributions of the attention and MLP blocks to the competition of the mechanisms. The attention blocks (left) contribute more to the marginal win of the counterfactual mechanism than the MLP blocks (right).
@@ Results.
Our results reveal the prevalence of each mechanism by varying the layer $l$ and position $i$.
Finding 1: Information flows from different tokens for different mechanisms. We analyze the role of previous context at various token positions with respect to different depths of the layers. In Figure 2(a), the blue heatmap above shows the logits of the factual token $t_{\mathrm{fact}}$, and the red heatmap below shows those of the counterfactual token $t_{\mathrm{cofa}}$.
Looking at the blue heatmap, we see that the subject position is the main contributor to the logits of $t_{\mathrm{fact}}$ in early layers, which is consistent with a previous finding Geva et al. (2023). Specifically, we also locate the factual attribute in the subject positions by the first MLP layer, and find they increase on average the value of $t_{\mathrm{fact}}$ from $0.38$ to $0.74$ in the premise and from $0.9$ to $1.93$ in the second sentence. Then, in the later layers, the strongest contributor is the last tokens before the attribute, as the last token position is used to predict the attribute. From the red heatmap, we see the evolution of $t_{\mathrm{cofa}}$’s logits. The observations of later layers are similar across two mechanisms, in that the last token contributes the most. However, in early layers, the counterfactual mechanism’s $t_{\mathrm{cofa}}$ token is best encoded in the attribute position instead of the subject position for the factual mechanism.
Such information flow between different token positions suggests a major role played by the attention mechanism in moving such information to the last position, resonating with observations in Geva et al. (2023).
Finding 2: Both the individual mechanisms and competition take place in late, but not early layers. We trace the competition of the two mechanisms across the layers by plotting in Figure 2(b) the scale of the logits corresponding to the two mechanisms in the last token position. The first observation is that the strength of each individual mechanism increases monotonically across the layers, from a relatively small logit below 1 in early layers to large values of around 15 in the final layer.
Another observation is that, although both mechanisms increase in strength, stronger signals of the competition (where the counterfactual mechanism prevails the factual one) start after the fifth layer, and this prevalence gradually grows in later layers. The logits of the counterfactual mechanism are, in most of the examples, the highest in the 50K-dimensional vocabulary of GPT-2, making $t_{\mathrm{fact}}$ dominant in 96% of the examples.
6.2 Intermediate Inspection of Attention and MLP Blocks
Behind the overall win of the counterfactual mechanism, we want to trace the contributions from the attention and MLP blocks in each layer.
Method.
For each attention or MLP block, it processes the input embedding and outputs the logits of $t_{\mathrm{fact}}$ and $t_{\mathrm{cofa}}$ to be added to the residual stream. We can consider the contribution of each block as its added logit values to the residual stream. Intuitively, if the added logit value for $t_{\mathrm{cofa}}$ is higher than that of $t_{\mathrm{fact}}$, then this block pushes the overall prediction to lean towards the counterfactual mechanism; otherwise, this block suppresses the counterfactual mechanism.